An object is thrown upward from a height of 80 ft. The initial velocity of the object is 64 ft per second. If the height h (in feet) is h=-16t²+64t+80, where t is time in seconds, when will the object reach the ground?
Answer by ikleyn(52915)

.
An object is thrown upward from a height of 80 ft. The initial velocity of the object is 64 ft per second.
If the height h (in feet) is h=-16t²+64t+80, where t is time in seconds, when will the object reach the ground?
~~~~~~~~~~~~~~~~

To find this time, solve the equation


    -16t^2 + 64t + 80 = 0,


saying that the height of the object is zero at the time t.


To facilitate the solution, first divide all the addends by -16. You will get


    t^2 - 4t + 5 = 0.


By the obvious way, it is factorable


    (t-5)*(t+1) = 0.


Of the two roots, t= 5 and t= -1, you choose the positive one and reject the negative one.


ANSWER.  In 5 seconds.

Solved.

-----------------

In this site,  there is a bunch of lessons on a projectile thrown/shot/launched vertically up
    - Introductory lesson on a projectile thrown-shot-launched vertically up
    - Problem on a projectile moving vertically up and down
    - Problem on an arrow shot vertically upward
    - Problem on a ball thrown vertically up from the top of a tower
    - Problem on a toy rocket launched vertically up from a tall platform

Consider these lessons as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.
Read them attentively and learn how to solve this type of problems once and for all.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Projectiles launched/thrown and moving vertically up and dawn".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.